OpenStudy (anonymous):

18n+70 15n+90 Will these two ever have equivalent answers? I know the answer is no but I need to know WHY.

OpenStudy (anonymous):

\[ 18n+70=15n+90 \]

OpenStudy (anonymous):

Solve for \(n\).

OpenStudy (anonymous):

There might not be any solution where \(n\in\mathbb N\).

OpenStudy (anonymous):

what i want to know is WHY they cant be equivalent

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

\[ 18n+70=15n+90\\ 3n = 20\\ n=\frac{20}{3} \]

OpenStudy (anonymous):

Two lines can only intersect at a single point.

OpenStudy (anonymous):

Unless they are parallel, then there will be 0 or infinite solutions.

OpenStudy (anonymous):

@wio I think I didn't give out enough info. Here's the full question: Is there a number of times for which A will be the same as B? Explain how you determined your answer.